A new project is expected to generate $1,000,000 in revenues, $250,000 in cash operating expenses, and depreciation expense of $200,000 in each year of its 10-year life. The corporation's tax rate is 21%. The project will require an increase in net working capital of $85,000 in the beginning and a decrease in net working capital of $75,000 in year ten. What is the free cash flow from the project in year one?

A) $298,000

B) $634,500

C) $380,000

D) $410,000
Answer #1

Revenues = $ 1000000, Cash Operating Expense = $ 250000, Depreciation Expense = $ 200000

Profit Before Tax (PBT) = Revenue - Cash Operating Expense - Depreciation Expense = 1000000 - 250000 - 200000 = $ 550000

Tax Expense @ 21 % = 0.21 x 550000 = $ 115500

Net Income = PBT - Tax Expense = 550000 - 115500 = $ 434500

Free Cash Flow in Year 1 = Net Income + Depreciation Expense = 434500 + 200000 = $ 634500

Hence, the correct option is (B)